OpenStudy (anonymous):

Attention future IITians. This is a tricky and witty question. Just try. Everyone should try. Prove that 1 = 2.

OpenStudy (mani_jha):

\[(1-2)^{2}=(2-1)^{2}=1\] \[1-2=2-1\] \[2=4\] \[1=2\] lol, we destroy mathematics!

OpenStudy (anonymous):

You didn't destroy it, you abused it.

OpenStudy (anonymous):

Some smart reply, Ishaaan Bhai, plzzzzzzzzzzz? Try some other method also.

OpenStudy (anonymous):

It's not possible adarsh, no matter what you do, you can't prove that

OpenStudy (anonymous):

Nothing is possible. Just try. I will post it soon.

OpenStudy (mani_jha):

Chill Ishaan, it was supposed to be a joke!

OpenStudy (anonymous):

Lol why do you think I am serious about it

OpenStudy (anonymous):

let x=x x^2=x^2 x^2-x^2=x^2-x^2 x(x-x)=(x+x)(x-x) x=2x therefore 1=2 though this proof does not mathematically exist as it is a falacy

OpenStudy (anonymous):

Really witty answer!!!!!!!

OpenStudy (anonymous):

All who can answer these type of questions can aim to get "Abel Prize".

OpenStudy (anonymous):

u can do 1=2 with binomial thm

OpenStudy (anonymous):

All of u r requested to check this link: http://www.resonance.ac.in/reso/downloads/FAQ_ISEET.pdf